OpenStudy (anonymous):

A line passes through the point (2, 3) and has a slope of -2. Which is the equation of the line in point-slope form?

OpenStudy (aripotta):

do you know what point-slope form is?

OpenStudy (anonymous):

i do not

OpenStudy (aripotta):

y - y1 = m(x - x1)

OpenStudy (anonymous):

oh ok ok

OpenStudy (aripotta):

y1 and x1 is where you'll substitute the x and y in the point given

OpenStudy (anonymous):

soo will it be y - 3 = -2(x - 2)

OpenStudy (aripotta):

yes :]

OpenStudy (anonymous):

thank you Ari

OpenStudy (aripotta):

no problemo~